OpenStudy (anonymous):

What are the possible rational zeros of f(x) = x4 - 4x3 + 9x2 + 5x + 14

OpenStudy (campbell_st):

the ration zero theorem says find the factors of the constant and let them be q find the factors of the coefficient of the leading term and let them be p then the possible rational zeros will be q/p so you are looking at \[\pm 1,\pm 2, \pm 7, \pm 14\]

OpenStudy (anonymous):

thank you @campbell_st can you answer this question as well? What are the possible number of positive real, negative real, and complex zeros of f(x) = 4x3 + x2 + 10x - 14

OpenStudy (anonymous):

i got 1,3,9

OpenStudy (anonymous):

plus or minus

OpenStudy (campbell_st):

well the factors of the constant are \[q = \pm 1, \pm 2, \pm 7, \pm 14\] and the leading term is \[p = \pm1, \pm2, \pm4\] so you need to now write them as \[\frac{q}{p} = \frac{\pm1}{\pm1}, \frac{\pm2}{\pm1}, \frac{\pm7}{\pm1} \frac{\pm14}{\pm1}\] and do the same with a denominator or 2 and repeat agaib with a denominator of 4 hope it helps

OpenStudy (anonymous):

thank you